OpenStudy (anonymous):

A bag contains n cards each having one of the numbers (1,2,3,4,...,n) written on it. If all numbers are used the probability of drawing a card with a number less than or equal to 10 is 4/10. How many cards are in the bag?

OpenStudy (anonymous):

The set should be {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, \ 20, 21, 22, 23, 24, 25} 10/25=40/100=4/10